The front 1.20 m of a 1,450-kg car is designed as a "crumple zone" that collapses to absorb the shock of a collision. (a) If a car traveling 26.0 m/s stops uniformly in 1.20 m, how long does the collision last? s (b) What is the magnitude of the average force on the car? N (c) What is the magnitude of the acceleration of the car? Express the acceleration as a multiple of the acceleration of gravity. g
Answer #1

a)

vi = initial speed = 26 m/s

vf = final speed = 0 m/s

d = stopping distance = 1.20 m

t = time taken = ?

Using the equation

d = (vf + vo) t/2

1.20 = (26 + 0 )t/2

t = 0.0923 sec

b)

magnitude of average force is given as

F = m (vf - vi)/t

F = (1450) (0 - 26)/0.0923

F = - 408450.7 N

magnitude = 408450.7 N

c)

acceleration is given as

a = (vf - vi)/t

a = (0 - 26)/0.0923

a = - 281.7 m/s2

magnitude : 281.7 = 281.7 g/9.8 = 28.7 g